OpenStudy (anonymous):

Solve. 38-7r=3(6-4r) Enter only the value of the variable.

OpenStudy (anonymous):

20=(-12+7)r r=20/(-5)=-4

OpenStudy (anonymous):

r = -4

OpenStudy (anonymous):